The mean yield of a particular variety of grain crop is 200 bu/acre (bushels per acre). If the top yielding 20% of the plants are selected for breeding and mated randomly among themselves, this is equivalent to a selection differential S of S = 37 bu/acre. What is the expected mean yield of the progeny of the selected parents if the heritability of yield is 0.428? (hint: don’t forget that the response (R) to selection represents the new value of a trait in the offspring generation, not the change in that value).
Answer #1

We have,

Selection Differential = 37 bu/acre

Selection Response

= Average yield of Progeny - Average yield of parental crop

= A - 200 bu/acre

h2 = 0.428

From the Breeder's Equation:

h2 = R/S

=> 0.428 = (A - 200)/37

=> A = (0.428 * 37) + 200 = 215.836

The means that the Mean Yield of the Progeny population is 215.836 bushels/acre